# Neurodynamics-Driven Coupled Neural P Systems for Multi-Focus Image Fusion

### v1

@@ 1 Introduction
In recent years, DL-based approaches have become the dominant paradigm in MFIF, achieving remarkable progress [16]. Some studies have applied DL techniques to end-to-end image fusion frameworks [34, 32]. Different end-to-end methods may use various network structures, but they generally consist of an encoder network for extracting image features, a fusion network for merging features, and a decoder network for reconstructing the fused features. Another line of studies draws inspiration from spatial domain methods, suggesting the use of DL techniques to generate decision maps [19, 28, 24, 42]. This process involves training a network to map the source image to a focus map, ultimately producing decision maps. Compared to end-to-end methods, decision-based methods directly retain information from the source image at the pixel level, avoiding the spatial consistency issues of end-to-end methods. Therefore, more research has been done on DL-based decision map methods. However, DL-based decision map methods still have evident drawbacks. The most crucial one is that the internal working mechanism is difficult to explain, hindering a clear understanding of the mapping between source images and the decision map [29]. This will result in nuisance false edges and burrs in the generated decision map, as can be intuitively observed in Fig. 1(a).
@@ 2.1 DL-based MFIF methods
End-to-end methods. These methods utilize the powerful fitting capabilities of deep neural networks to directly generate fused images by learning features from the source images. Representative approaches include those based on convolutional neural networks [41, 31, 35, 34, 3], generative adversarial networks [36, 6], and Transformer architectures [20, 43, 26]. It is challenging for these methods to maintain the spatial consistency of the source images.
@@ 3.2 Why neurodynamics analysis is necessary?
Figure 2: CNP neuron firing states on different inputs. (a) $I=0.5$; (b) $I=1.0$; (c) $I=1.5$; and (d) $I=2.0$.
@@
Continuous firing requires the fundamental criterion $U(t)\times(1+V(t))>T(t)$ to hold at all times. Substituting results $U(t)$, $V(t)$ and $T(t)$ obtained by Theorems 12 and 3, we can obtain:
$\begin{split}&\left(I\frac{1-\alpha^{t}}{1-\alpha}+\sum_{i=0}^{t-1}K(i)\alpha^{t-i-1}\right)\left(1+\sum_{i=0}^{t-1}K(i)\beta^{t-i-1}\right)\\ &\quad\quad\quad\quad\quad\quad\quad\quad\quad>\lambda\frac{1-\gamma^{t-1}}{1-\gamma}.\end{split}$
@@
When $t\to\infty$, we obtain Eq. (19)
@@
$I<\frac{\lambda(1-\alpha)(1-\beta)}{(1-\gamma)(1-\beta+\operatorname{sum}(W)))}-\operatorname{sum}(W).$ (21)
Corollary 1 reveals the necessary constraints between parameters and input signals for the proper functioning of CNP neurons. Detailed CNP parameter configurations (e.g., $\tau$ and $\lambda$) derived from Corollary 1 are provided in Appendix A.2. Noting that parameters are automatically configured by the system based on input MFIF images, eliminating the manual tuning and demonstrating transferability and generalization across different datasets. This helps to improve the ability of CNP systems to distinguish between focused and defocused regions in MFIF tasks, thereby improving the quality of the generated decision map.
@@ 4.2 Overview of the architecture
Table 2: Runtime comparison of different fusion methods. Note that PADCDTNP, SAMF, and our proposed method are executed on an i5-13400 CPU, while the remaining methods are executed on an A100 GPU.
@@ 5.2 Performance comparison
Runtime comparison. Tab. 2 lists the running times of all fusion methods on the Lytro dataset. Our method ranks third in runtime, but can achieve better performance within an acceptable execution time. Our ND-CNPFuse implementation is currently in MATLAB, and a re-implementation in C++ would further improve efficiency.
@@ 5.4 Ablation study
Neurodynamic analysis. We conducted ablation studies on the Lytro dataset to validate the effectiveness of neurodynamic analysis in improving model performance. Specifically, we compare the baseline CNP system and the neurodynamics-driven CNP system across six metrics, as summarized in Tab. 3. We can find that the neurodynamics-driven CNP system achieves significant improvements over the baseline across six multidimensional image quality metrics, with improvements of 2.02%, 17.09%, 1.56%, 5.01%, 1.45%, and 7.28%. These findings confirm the effectiveness of neurodynamic analysis.
SML module. We further validated the impact of the SML module on overall architecture performance. According to the quantitative analysis results in Tab. 4, after removing the SML module, evaluation metrics showed only a slight downward trend (e.g., PSNR decreased by 0.03 dB, and SSIM decreased by less than 0.02). This confirms the generalizability of the proposed method. Of course, by applying SML, the quality of fusion can be further enhanced.
We believe that ND-CNPFuse is unrestricted in the level of generalizability. First, neurons all parameters are automatically configured via neurodynamics. Second, the introduction or removal of the SML module has no significant effect on the overall framework’s performance.
@@ 5.5 Case study
We perform a qualitative case study on some samples from the Lytro dataset. Fig. 8 shows decision maps from the baseline and ND-CNP systems. The ND-CNP system produces decision maps with clearer boundaries and higher accuracy. In contrast, the original CNP system has uncontrolled spike firing in some areas, resulting in obvious misjudgments in the decision maps.
@@ 6 Conclusion
In this paper, we propose a novel MFIF method called ND-CNPFuse based on neurodynamics-driven CNP systems. The main contributions of ND-CNPFuse have two aspects. The first is that we provide a detailed theoretical neurodynamic analysis of the working mechanism of CNP neurons. This analysis reveals the constraints between network parameters and input signals, which helps prevent model failure caused by continuous firing. The second one is that we use spike counts of neurons to distinguish between focused and unfocused regions to generate decision maps without any post-processing, and the process is interpretable. Extensive experiments with nine SOTA fusion methods on four datasets show that our method is more competitive and generalizable.

### v2

@@ 1 Introduction
In recent years, DL-based approaches have become the dominant paradigm in MFIF, achieving remarkable progress [16]. Some studies have applied DL techniques to end-to-end image fusion frameworks [35, 33, 26]. Different end-to-end methods may use various network structures, but they generally consist of an encoder network for extracting image features, a fusion network for merging features, and a decoder network for reconstructing the fused features. Another line of studies draws inspiration from spatial domain methods, suggesting the use of DL techniques to generate decision maps [19, 29, 24, 43, 7]. This process involves training a network to map the source image to a focus map, ultimately producing decision maps. Compared to end-to-end methods, decision-based methods directly retain information from the source image at the pixel level, avoiding the spatial consistency issues of end-to-end methods. Therefore, more research has been done on DL-based decision map methods. However, DL-based decision map methods still have evident drawbacks. The most crucial one is that the internal working mechanism is difficult to explain, hindering a clear understanding of the mapping between source images and the decision map [30]. This will result in nuisance false edges and burrs in the generated decision map, as can be intuitively observed in Fig. 1(a).
@@ 2.1 DL-based MFIF methods
End-to-end methods. These methods utilize the powerful fitting capabilities of deep neural networks to directly generate fused images by learning features from the source images. Representative approaches include those based on convolutional neural networks [42, 32, 36, 3], generative adversarial networks [37, 6], and Transformer architectures [20, 44, 27]. It is challenging for these methods to maintain the spatial consistency of the source images.
@@ 3.2 Why neurodynamics analysis is necessary?
Figure 2: CNP neuron firing states under different inputs. (a) $I=0.5$; (b) $I=1.0$; (c) $I=1.5$; and (d) $I=2.0$. When II exceeds the continuous firing condition, the neuron exhibits an abnormal state, as observed in (c) and (d).
@@
Continuous firing requires the fundamental criterion $U(t)\cdot(1+V(t))>T(t)$ to hold at all times. Substituting results $U(t)$, $V(t)$ and $T(t)$ obtained by Theorems 12 and 3, we can obtain:
$\begin{split}&\biggl(I\frac{1-\alpha^{t}}{1-\alpha}+\sum_{i=0}^{t-1}K(i)\alpha^{t-i-1}\biggl)\biggl(1+\sum_{i=0}^{t-1}K(i)\beta^{t-i-1}\biggl)\\ &\quad\quad\quad\quad\quad\quad\quad\quad>\lambda\frac{1-\gamma^{t-1}}{1-\gamma}.\end{split}$ (17)
@@
When $t$ goes to infinity, we obtain Eq. (19)
@@
$I\leq\frac{\lambda(1-\alpha)(1-\beta)}{(1-\gamma)(1-\beta+\operatorname{sum}(W)))}-\operatorname{sum}(W).$ (21)
Corollary 1 reveals the necessary constraints between parameters and input signals for the proper functioning of CNP neurons. Detailed CNP neurons parameter configurations (e.g., $u$, $v$, $\tau$, $\lambda$, and W) derived from Corollary 1 are provided in Appendix A.2. Noting that the internal parameters of neurons are automatically configured based on input MFIF images, eliminating the manual tuning, and demonstrating transferability and generalization across different datasets. This helps to improve the ability of CNP systems to distinguish between focused and defocused regions in MFIF tasks, thereby improving the quality of the generated decision map.
@@ 4.2 Overview of the architecture
Table 2: Runtime comparison of different fusion methods. PADCDTNP, SAMF, and our method are executed on an i5-13400 CPU, while the remaining methods run on an A100 GPU. Our method reports dual runtimes: 0.41s (MATLAB) and 0.18s (C++) per image pair.
@@ 5.2 Performance comparison
Runtime comparison. Tab. 2 lists the running times of all fusion methods on Lytro. Our MATLAB implementation achieves 0.41s per image pair. By migrating to C++, the runtime is reduced to 0.18s, showing even GPU-accelerated DMANet (0.21s) in CPU-only execution. This confirms that our approach meets real-time MFIF requirements.
@@ 5.3 Parameter sensitivity analysis
In summary, the consistent performance across the four datasets demonstrates the universality of $t$ and $r$.
@@ 5.4 Ablation study
Neurodynamic analysis. We conducted ablation studies on the Lytro dataset to validate the effectiveness of neurodynamic analysis in improving model performance. Specifically, we compare the baseline CNP system and the neurodynamics-driven CNP system across six metrics, as summarized in Tab. 3. We can find that the neurodynamics-driven CNP system achieves significant improvements over the baseline across six multidimensional image quality metrics, with improvements of 2.01%, 17.29%, 1.55%, 5.01%, 1.44%, and 7.29%. These findings confirm the effectiveness of neurodynamic analysis.
SML module. We further validated the impact of the SML module on overall architecture performance. According to the quantitative analysis results in Tab. 4, after removing the SML module, evaluation metrics showed only a slight downward trend (e.g., PSNR decreased by 0.007 dB, and SSIM decreased by 0.002). This confirms the generalizability of the proposed method. Of course, by applying SML, the quality of fusion can be further enhanced.
We believe that ND-CNPFuse is unrestricted in the level of generalizability. First, all internal parameters of neurons are automatically configured via neurodynamics. Second, the introduction or removal of the SML module has no significant effect on the overall performance of the framework.
@@ 5.5 Case study
We perform a case study on the Lytro dataset samples. Fig. 8 shows decision maps from the baseline and ND-CNP systems. The ND-CNP system produces decision maps with clearer boundaries and higher accuracy. In contrast, the original CNP system has uncontrolled spike firing in some areas, resulting in obvious misjudgments in decision maps.
@@ 6 Conclusion
In this paper, we propose a novel MFIF method called ND-CNPFuse based on neurodynamics-driven CNP systems. The main contributions of ND-CNPFuse have two aspects. The first is that we provide a detailed theoretical neurodynamic analysis of the working mechanism of CNP neurons. This analysis reveals the constraints between network parameters and input signals, which helps prevent model failure caused by continuous firing. The second one is that we use spike counts of neurons to distinguish between focused and unfocused regions to generate decision maps without any post-processing, and the process is interpretable. Extensive experiments with nine SOTA fusion methods on four datasets show that our method is more competitive.